# CMMCoT: Enhancing Complex Multi-Image Comprehension via Multi-Modal Chain-of-Thought and Memory Augmentation

### v1

@@ 4.4 Performance Evaluation
Multi-image performance: The experimental results across multiple diagnostic benchmarks demonstrate the consistent superiority of our model across both 7B and 2B parameter configurations. As shown in Table  1, our 7B variant achieves state-of-the-art performance on Mantis-Eval (70.9) and competitive results on NLVR2 (88.7) and BLINK (52.3), outperforming comparable 7B models like qwen2-vl by 2.6 average points. Notably, the scaled-down 2B version maintains robust capabilities, surpassing its parameter-matched counterpart qwen2-vl-2B by 2.4 average points while exhibiting particular strengths in temporal reasoning (MVBench: 40.1) and visual-textual verification (NLVR2: 76.9). This dual-scale effectiveness suggests our architecture successfully balances model capacity with cross-modal alignment efficiency, preserving critical visual reasoning abilities even in parameter-constrained scenarios. The performance parity across distinct evaluation paradigms from multi-image logical analysis (Mantis) to perceptual similarity judgments (BLINK), validates our approach’s adaptability in handling diverse visual reasoning tasks through enhanced feature interaction mechanisms.
@@
Single-image performance: Similarly, as shown in Table 2 in single-image datasets, our model consistently achieves superior performance across different scales compared to prior architectures, demonstrating that the proposed multi-modal hybrid sequential chain reasoning strategy effectively enhances comprehension capabilities for single-image tasks. This cross-scale effectiveness confirms the generalizability of our methodology beyond multi-image analytical scenarios, particularly in improving fundamental visual-semantic alignment through structured reasoning pathways.
4.5 Ablation Study
We also conducted a series of ablation experiments to investigate the impact of distinct components and features on model performance, as shown in Table 5. First, we analyzed the scenario where only grounding-based information was utilized for training. The experimental results demonstrated that the exclusive incorporation of grounding information yielded marginal performance improvements, specifically manifesting as a 0.2-point increase in the average score. However, this enhancement, though observable, was not statistically significant. Subsequently, we explored the synergistic effects of integrating both grounding information and entity images during the training phase. The outcomes proved highly encouraging, with the model’s average score substantially improving from 64.5 to 66.3. This significant gain indicates that the combination of grounding information with visual entity features effectively enhances the model’s capacity to comprehend and process multimodal representations.
Furthermore, we extended our investigation by implementing the RIFREM module during the inference stage. Empirical evidence revealed that this strategy delivered additional incremental performance gains, elevating the average score from 66.3 to 67.1. These results validate the efficacy of the RIFREM module in excavating and leveraging cross-image feature relationships, thereby contributing to enhanced overall performance.
Collectively, these systematic experiments not only delineate the individual contributions of each component but also reveal the latent potential of coordinated feature utilization in multimodal architectures. Our findings suggest that practical implementations should holistically consider these synergistic mechanisms to optimize model performance and application adaptability.
@@ 4.5 Ablation Study
Table 3: Performance Comparison of Qwen2-VL Variants Across Multiple Benchmarks
Models BLINK Mantis NLVR2 MVBench Q-Bench Avg
qwen2-vl $51.8$ $64.8$ $84.0$ $50.5$ $71.2$ $64.5$
qwen2-vl-grounding $51.6$ $65.4$ $85.8$ $51.1$ $69.9$ $64.7$
qwen2-vl-cmmcot $52.6$ $66.8$ $88.5$ $50.0$ $73.4$ $66.3$
qwen2-vl-cmmcot-RIFREM $52.3$ $70.9$ $88.7$ $51.3$ $72.2$ $67.1$
4.6 Analysis of RIFREM
Incorporating the RIFREM module during inference enables the mining of multi-graph features associated with target entities. However, this module introduces additional inference latency. We systematically investigated the impact of integrating RIFREM at different network depths on both performance and latency through controlled experiments, as shown in Figure 4. Five architectural configurations were evaluated: Group 1 (layers 0 and 27), Group 2 ([0, 9, 18, 27]), Group 3 ([0, 4, 8, 12, 16, 20, 24, 27]), Group 4 ([0, 1, 3, 5, 7, 9, 11, 13, 14, 16, 18, 20, 22, 24, 26, 27]), and Group 5 (all layers).
Experimental results demonstrate that restricting RIFREM integration to only the initial and final layers (Group 1) leads to performance degradation. The underlying reason may be attributed to the disruption of information propagation and consistency preservation across different network depths when applying feature enhancement solely at terminal layers. Progressive integration through multiple depth levels (as in Groups 2-5) facilitates iterative feature refinement and consistency maintenance throughout the network hierarchy. To achieve optimal balance between model effectiveness and computational efficiency, we adopted the Group 3 configuration with 8 strategically distributed layers. All subsequent experimental analyses were conducted under this optimized architecture setting.
4.7 Quantitive Analysis
To further investigate how multimodal hybrid sequence Chain-of-Thought (CoT) enhances model comprehension of complex multi-image scenarios, we visualize several representative cases for detailed analysis, as demonstrated in Figures 1 and 5.
In the case shown in Figure 1, comparative analysis reveals that our model outperforms both Qwen2-VL and GPT-4O by successfully extracting rock regions from each image and conducting structural stability comparisons across different geological formations. Our approach provides highly convincing visual evidence and textual rationale, demonstrating enhanced spatial reasoning capabilities through systematic region segmentation and comparative analysis.
The left demo of Figure 5 illustrates our model’s superior performance in object detection and quantitative reasoning. While competing models produced erroneous answers, our architecture achieves precise localization and recognition of cutlery items (knives and spoons), followed by accurate quantity enumeration and comparative judgment. This systematic process highlights our method’s improved counting capability, particularly in resolving ambiguous object instances that challenge baseline models.
As shown in the right demo of Figure 5, although alternative models can produce correct answers, our framework demonstrates advanced logical reasoning through anomaly detection. Specifically, our model identifies and visualizes regions violating commonsense logic within the visual context, accompanied by comprehensive causal analysis. This capability to detect and explain contextual inconsistencies confirms the enhanced logical reasoning capacity of our approach.
5 Conclusion
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A You can have an appendix here.
Table 4: The statistics of MMCoT-260K
@@
Table 5: Statistics of the evaluation dataset. For MVBench, we extract 8 frames per video for the evaluation.
Benchmarks Skill Type Number Avg-I Max-I
BLINK Reason MCQ $1901$ $1.9$ $4.0$
Mantis Reason/Coref MCQ/Short $217$ $2.5$ $5.0$
NLVR2 Reason MCQ $6967$ $2.0$ $2.0$
MVBench Temporal MCQ $4000$ $8.0$ $8.0$
Q-Bench Comparison MCQ $1000$ $2.0$ $2.0$
We also provide additional visualization examples, as shown in Figure 6 7, 8, including a bad case 9. These successful cases further demonstrate the efficacy of our model. From the bad case, we can observe that despite the model making an incorrect prediction, it is still capable of extracting and analyzing key visual elements mentioned in the text.
@@
Figure 8: More examples of multi-image tasks.
Figure 9: Bad cases in complex multi-image tasks.

### v2

@@ Methods
Multi-image performance: The experimental results across multiple diagnostic benchmarks demonstrate the consistent superiority of our model across both 7B and 3B parameter configurations. As shown in Tab. 1, our 7B variant achieves state-of-the-art performance on Mantis-Eval (72.2), MVBench(75.8), Q-Bench(78.5), and competitive results on NLVR2 (89.9) and BLINK (52.3), outperforming comparable 7B models like Qwen2.5-VL by 1.4 average points. Notably, the scaled-down 3B version maintains robust capabilities, surpassing its parameter-matched counterpart Qwen2.5-VL-3B by 2.6 average points while exhibiting particular strengths in temporal reasoning (MVBench: 73.1) and visual-textual verification (NLVR2: 88.9). This dual-scale effectiveness suggests our architecture successfully balances model capacity with cross-modal alignment efficiency, preserving critical visual reasoning abilities even in parameter-constrained scenarios. The performance parity across distinct evaluation paradigms from multi-image logical analysis (Mantis) to perceptual similarity judgments (BLINK) validates our approach’s adaptability in handling diverse visual reasoning tasks through enhanced feature interaction mechanisms.
@@
Single-image performance: Similarly, as shown in Tab. 2 in single-image datasets, our model consistently achieves superior performance across different scales compared to prior architectures, demonstrating that the proposed multi-modal hybrid sequential chain reasoning strategy effectively enhances comprehension capabilities for single-image tasks. This cross-scale effectiveness confirms the generalizability of our methodology beyond multi-image analytical scenarios, particularly in improving fundamental visual-semantic alignment through structured reasoning pathways.
Quantitive Analysis
To further evaluate how multimodal hybrid sequence CoT enhances model understanding in complex multi-image scenarios, we visualize representative examples in Fig. 3. The left example in Fig. 3 highlights our model’s superiority in object localization and quantitative reasoning. While other models fail to correctly detect or count cutlery items (e.g., knives and spoons), our model accurately identifies their positions and quantities, demonstrating robust counting ability even in ambiguous visual contexts. The right example illustrates our model’s advantage in anomaly detection and logical reasoning. Although alternative models may provide correct answers, our method not only identifies commonsense violations in the visual input but also offers explicit causal interpretations. This capacity for contextual inconsistency detection underscores the enhanced reasoning capabilities enabled by our approach.
Ablation Study
Impact of Different Modules. We conducted a series of ablation studies to examine the contributions of individual components, as shown in Table 3. We first analyzed the effect of using only grounding-based information during training. This setting yielded a modest 0.1-point improvement in average score, suggesting limited benefit when used in isolation. Next, we investigated the impact of combining grounding with entity images. This joint supervision led to a notable performance gain, improving the average score from 73.2 to 74.4. The result highlights the synergy between grounding cues and visual features in enhancing multimodal representation learning. We further evaluated the RIFREM module applied during inference. Its integration provided an additional boost, raising the average score to 74.6 by exploiting cross-image feature relationships. Together, these experiments clarify the individual and combined contributions of each module. They underscore the importance of coordinated feature utilization—across both training and inference—for optimizing performance in multimodal tasks.
Impacts of RIFREM. We evaluate the effect of integrating the RIFREM module at different network depths during inference, balancing performance gains and latency overhead (Figure 4). Five configurations were tested, from shallow (Group 1: layers 0 and 27) to full-depth integration (Group 5: all layers). Results show that applying RIFREM only at the first and last layers degrades performance, likely due to disrupted information flow. In contrast, progressive integration across the network (Groups 2–5) enables iterative refinement and better feature consistency. Group 3 (8 evenly distributed layers) achieves the best trade-off between accuracy and efficiency, and is used in all subsequent experiments.
@@
Conclusion
@@
Acknowledgments
This research was supported in part by Zhejiang Provincial Natural Science Foundation of China under Grant LD24F020016, the Key R&D Program of Zhejiang Province 2025C01075, 2023C01043, the National Natural Science Foundation of China under Grant 62576313, and Alibaba Group through Alibaba Research Intern Program.
Appendix
More Experimental Results
Fine-tuned results of different models
To further verify that the innovation of our approach is not solely dependent on the dataset, we analyzed the results of different models fine-tuned on the CMMCoT-260k dataset. As shown in 4, the experimental results show that the LLaVA and Mantis models exhibit a significant performance decline after CMMCoT fine-tuning, likely due to their lack of visual localization capabilities, which is exacerbated by the fine-tuning process. The LLaVA-OV model, which incorporates localization data during pre-training, shows a smaller degree of performance degradation, yet still experiences some decline, possibly due to the absence of explicit localization mechanisms in its output. The Qwen2-VL model, equipped with localization capabilities through pre-training and architectural design, demonstrates moderate improvement after fine-tuning. The results confirm that only when the original multimodal model possesses strong localization abilities can the full potential of the CMMCoT method be realized.
Table 4: Performance comparison across multimodal benchmarks.
Model Training BLINK Mantis NLVR2 MVBench Q-Bench Avg
LLaVA-v1.5 (7B) Frozen 37.1 41.9 52.1 36.0 53.9 44.2
LLaVA-v1.5 (7B) Fine-tuned 35.5 30.4 49.0 32.7 40.2 37.6
LLaVA-v1.6 (7B) Frozen 39.6 45.6 58.9 40.9 58.9 48.8
LLaVA-v1.6 (7B) Fine-tuned 38.5 40.8 53.2 38.2 55.9 45.3
Mantis-Idefics2 (8B) Frozen 49.1 57.1 89.7 51.4 75.3 64.5
Mantis-Idefics2 (8B) Fine-tuned 46.3 59.4 84.6 40.8 69.6 60.1
LLaVA-OV (7B) Frozen 48.2 64.2 89.4 56.7 74.5 66.6
LLaVA-OV (7B) Fine-tuned 46.9 62.5 87.4 50.8 73.3 64.2
Qwen2-VL (7B) Frozen 51.8 64.8 84.0 50.5 71.2 64.5
Qwen2-VL (7B) Fine-tuned 52.6 66.8 88.5 50.0 73.4 66.3
Qwen2.5-VL (7B) Frozen 55.3 69.8 88.3 74.7 77.7 73.2
Qwen2.5-VL (7B) Fine-tuned 55.6 70.6 88.7 74.9 77.9 73.5
Qwen2.5-VL(7B) Our Method 56.8 72.2 89.9 75.8 78.5 74.6
Details of CMMCoT-260K
CMMCoT-260K encompasses four distinct types of data: Caption, Co-reference, Comparison, and Reason.
Caption
The caption data type is employed to generate natural language descriptions for images, serving as a direct expression of the semantic information contained within the image. It provides the model with the ability to comprehend and semantically parse image content, facilitating tasks such as annotating and generating descriptions for multiple images.
Co-reference
Co-reference data involves identifying and linking elements that refer to the same entity across different images. This data type assists the model in understanding entity consistency within a single image or among multiple images. It enhances the model’s ability to maintain consistent entity reasoning across multiple images, thereby strengthening its understanding of complex scenes.
Comparison
The comparison data type is utilized for conducting comparative analyses between multiple images, identifying similarities and differences among elements in different images. It improves the model’s comparative and contrastive analytical capabilities, enabling it to make more nuanced and accurate judgments in tasks involving multiple images, such as in visual question answering.
Reason
The reason data type pertains to more complex reasoning tasks, requiring the model to perform logical reasoning and decision-making based on relationships among multiple images. This typically necessitates contextual understanding and causal relationship analysis between images. It enhances the model’s capacity to solve advanced reasoning problems, making it applicable for understanding more complex task contexts.
Table 5: The statistics of CMMCoT-260K
@@ Acknowledgments
More Examples
We also provide additional visualization examples, as shown in Figure 5, Figure 6 and Figure 7. These successful cases further demonstrate the efficacy of our model. From the bad case in Figure 7, we can observe that despite the model making an incorrect prediction, it is still capable of extracting and analyzing key visual elements mentioned in the text.
Figure 5: Visualization results of our CMMCoT task to illustrate the fine-grained reasoning ability of our model.